### v2

# SMMAN: quasi-Simultaneous Multi-wavelength Monitoring of gamma-ray-loud AGNs with the Nanshan 26-m radio telescope

Abstract
Active Galactic Nuclei (AGNs) are characterized by strong temporal flux density variability across the electromagnetic spectrum, offering insights into the complex physical processes governing accretion and plasma outflows. To systematically investigate AGNs flux density variability in radio bands, a long-term program was initiated in late 2016: quasi-Simultaneous Multiwavelength Monitoring of gamma-ray-loud AGNs with the Nanshan 26-m radio telescope (SMMAN). This work presents the first data release of the SMMAN program, spanning over eight years from 2016 to 2024 with observations at 4.8 and 23.6 GHz. The SMMAN sample includes 131 northern () sources selected from the Fermi Large Area Telescope third source catalog. The characteristics of variability, spectral index, luminosity, and -ray loudness factor are examined for different AGN classes within the sample. Target sources exhibit stronger variability at 23.6 GHz compared to 4.8 GHz, with BL Lac objects being more variable than flat-spectrum radio quasars (FSRQs). BL Lacs generally have flatter radio spectra, while FSRQs, blazar candidates of uncertain type (BCUs), and radio galaxies (RDGs) span a wider range from flat to steep. FSRQs are more radio-luminous than BL Lacs and other classes, with BCUs intermediate and RDGs generally fainter. FSRQs and BL Lacs have higher -ray loudness factors than RDGs, while BCUs have intermediate values. The SMMAN dataset, incorporated with other historical and ongoing monitoring programs, will provide a unique opportunity to investigate the evolution of spectral energy distributions, search for quasi-periodic oscillations, and analyze supermassive black hole binary systems.
I Introduction
Galaxies that contain a highly bright, compact region at their center, called active galactic nuclei (AGNs), are known as active galaxies (Urry and Padovani, 1995). It is found that the high luminosity at the central compact region is powered by supermassive black holes (SMBH; mass ) via processes known as accretion (Frank et al., 2002). Active galaxies are also found to possess relativistic jet structures emanating from the central region (Rees, 1966).
Active galaxy’s multiwavelength emission is dominated by non-thermal radiation
extending from radio to -ray bands exhibiting intense and rapid flux density variations. In particular, a class of AGNs called blazars with a relativistic jet pointing towards Earth show extreme variability (Blandford and Rees, 1978; Blandford and Königl, 1979).
Due to the small jet viewing angle, the blazar’s observed emission is highly Doppler boosted and outshines all other components, thus making them ideal candidates to study the physical conditions and emission processes in relativistic jets.
From an observational standpoint, blazars are categorized into two sub-classes: flat-spectrum radio quasars (FSRQs) and BL Lacertae (BL Lacs) objects. FSRQs exhibit emission line equivalent widths (EW) > Å, while BL Lacs show EW < Å (Stocke et al., 1991; Stickel et al., 1991). As per the classification suggested by Ghisellini et al. (2011), which relies on the luminosity of the broad-line region (BLR), FSRQs exhibit a higher BLR luminosity () compared to BL Lacs, with representing the Eddington luminosity of the supermassive black hole located at the core of the AGN. The flux variability timescales in blazars range from minutes to years (e.g., Boettcher, 2004; Foschini et al., 2013; Orienti et al., 2014; Böttcher, 2019, and references therein) and were classified into three types caused by different mechanisms (Fan, 2005). These variations occur as flares or the quiescent/low-activity states. Another key feature of active galaxies is that their broadband spectral energy distribution (SED) from radio to -rays exhibits two-hump structure (e.g. Fossati et al., 1998). In addition, it is observed that AGNs/blazars exhibit high optical and radio polarization during flaring periods along with their polarization, showing intensive variability not only in terms of fraction but also in terms of the orientation of the polarization plane (e.g. Jorstad et al., 2007; Marscher et al., 2010; Blinov et al., 2021).
Studying the physical properties of AGNs through multi-frequency data has been found to be one of the most effective ways in the past, which has been employed in several studies (Marscher 2016; Böttcher 2019, and the references therein).
A key feature of AGNs is that the flux density observed from them shows variations across different wavebands over varying time intervals. Further, it is observed that the behavior of the broadband
flares of blazars is quite complex. The literature study points out that such flaring
activities are not always simultaneous in different wavebands. Also, it is important to study the long-term activity of blazars. From the multiband observations of the AGNs, it is noticed that the variability patterns in different wavebands are sometimes correlated and sometimes not (e.g. Liodakis et al. 2018, and references therein). Comprehending the physical basis for the observed variability remains one of the key unresolved issues in the study of AGNs. Temporal studies can be an efficient method to extract the time evolution trends of observed variability and to identify plausible connections across various blazar emission bands (e.g. Liodakis et al. 2018, and references therein).
Therefore, a highly effective approach to comprehending jet emission and its dynamics involves examining the alterations in the physical characteristics of AGNs/blazars that are necessary to generate variations in multi-frequency light curves. Hence, simultaneous monitoring of the AGNs across different wavelengths is important to discern the physical mechanisms in them.
The “Fermi Gamma-ray Space Telescope” (hereafter Fermi) was launched in
June of 2008 and the “Large Area Telescope” (LAT; Atwood et al. 2009) on-board provides a unique opportunity for the systematic study of AGNs/blazar variability (Böttcher, 2019; Principe, 2022; Casaburo et al., 2025). Fermi-LAT with its initial four years of observation, the source catalog (3FGL; Acero et al. 2015) includes blazars that account for more than of the identified sources. The most recent update of the fourth source catalog (4FGL-DR4) spans more than 14 years and includes a total of 7,195 sources within the energy range of 50 MeV to 1 TeV (Abdollahi et al., 2020; Ballet et al., 2023). Approximately 56% of these sources are associated with AGNs, including blazars. Conducting a comprehensive monitoring campaign of AGNs across multiple bands (radio to -ray) and carrying out temporal and spectral analysis is essential to understand the causes of the underlying flux density variations in them. Under this motivation, prior to the launch of the Fermi and after it, there are a few ongoing and historical AGNs multiband monitoring programs111https://www.cv.nrao.edu/MOJAVE/blazarlist.html.
A few of the major radio band surveys are briefly introduced below, with emphasis on the importance of multi-frequency radio monitoring projects. The University of Michigan Radio Astronomy Observatory’s (UMRAO; Aller et al. 1985) 26-m paraboloid telescope was devoted to monitoring the flux density and polarization of blazar jets from the mid-1960s until June 2012 (Aller et al., 2017). This facility provided centimeter-band variability over decades at 14.5, 8.0, and 4.8 GHz. The UMRAO data have been used in numerous extensive studies concerning flaring blazars and to acquire information regarding various jet and shock parameters that are not directly observable (Aller et al., 2014; Hughes et al., 2015; Aller et al., 2016) and also in several multiband temporal and spectral studies (e.g. Krishna Mohana et al., 2024; Mohana A et al., 2025). UMRAO observations were also used for the calibration and interpretation of the less frequently sampled Very Long Baseline Array (VLBA) imaging data (Aller et al. 2017, and references therein). It is to be noted that the long-term monitoring data set could help to search for a supermassive black hole binary (SMBHB)
candidates by identifying the periodic nature in the light curves (e.g. Readhead et al., 2025; Molina et al., 2025).
The radio telescope RATAN-600 (Parijskij, 1993) in Zelenchukskaya, Russia, which is an ongoing survey, has regularly monitored AGNs, including blazars, since 2005.
The observations were carried out with the six frequency bands at 1.25, 2.25, 4.7, 8.2, 11.2, and 22.3 GHz, simultaneously (within 1–2 min) when a source moves along the focal line where the receivers are located. These data formed the basis for the first version of the BLcat (Mingaliev et al., 2014), and the latest updated version of this catalog is “RATAN-600 multi-frequency catalog of blazars” (Sotnikova et al., 2022), which includes about 1830 source. Using these catalogs with information of flux densities, spectral indices, variability, and radio luminosity for the large blazar list, several studies to understand the flare activity and synchrotron radio spectra of AGNs were reported (e.g. Mufakharov et al., 2015; Sotnikova et al., 2019; Larionov et al., 2020; Sotnikova et al., 2021; Krishna Mohana et al., 2024; Kunert-Bajraszewska et al., 2025). The Fermi-GST AGN Multi-frequency Monitoring Alliance (F-GAMMA) programme (Fuhrmann et al., 2007; Angelakis et al., 2010; Fuhrmann et al., 2014), conducted at the Effelsberg telescope and IRAM 30 m telescope in Germany, performed nearly monthly broadband radio monitoring of selected blazars from January 2007 to January 2015 (Angelakis et al., 2019a). The monitoring was carried out at multi-frequency from 2.64 to 43 GHz, which includes the frequencies of 2.64, 4.85, 8.35, 10.45, 14.6, 23.05, 32, 43, 86.2, 142.3, and 228.9 GHz. Using initial 3.5 yr data, a cross-correlation analysis was carried out by Fuhrmann et al. (2014) between the radio and -ray light curves of 54 Fermi-bright blazars, with the aim of determining the exact location of the -ray emission region. Fuhrmann et al. (2016) reported a multi-frequency study of AGNs in Fermi era to examine whether the -ray loudness is related to the radio variability. The 2007-2015 multi-frequency radio data published in Angelakis et al. (2019a) is used in several temporal and spectral studies of blazars (e.g. Paliya et al., 2021; Kim et al., 2022; Krishna Mohana et al., 2024).
The Owens Valley Radio Observatory (OVRO) blazar monitoring program (Richards et al., 2011) offers an extensive data collection to examine AGNs variability at 15 GHz. It stands as the most comprehensive and sensitive radio monitoring survey of blazars, which has been in progress since 2008. Currently, the complete sample encompasses approximately 1850 sources, each of which is observed at a frequency of roughly twice a week. Observations from OVRO helped make progress in blazar jet physics through cross-correlations in the time domain between Fermi-LAT -ray and 15 GHz light curves. It was found that there is a close connection between -ray and 15 GHz blazar emission (Max-Moerbeck et al., 2014; Hovatta et al., 2015; Pushkarev et al., 2019).
In 2004, a program was initiated to monitor blazars at the 32 m radio antennas of the INAF-Institute of Radioastronomy (IRA), located in Medicina and Noto, Italy (Bach et al., 2007).
Beginning in 2004, 47 targets were observed with a monthly cadence at 5, 8, 24, and 43 GHz.
The Radio Observations of Blazars with INaf telescopes (ROBIN) program has continued to thrive over the years and remains active at frequencies of 8 and 24 GHz, having reached 20 years of operation (Marchili et al., 2025a). The data accumulated through the program have been used in numerous studies and publications, primarily in relation to Whole Earth Blazar Telescope (WEBT) multi-frequency campaigns (e.g. Raiteri et al., 2007; Abdo et al., 2010b; Carnerero et al., 2015; Raiteri et al., 2017, 2024).
The monitoring program has facilitated the examination of how the variability characteristics and SED of blazars evolve over time (Marchili et al. 2025a, and references therein).
At the Aalto University Metsähovi Radio Observatory (MRO, Finland), several hundred AGNs and a few galactic microquasars have been monitored at 22, 37, and 87 GHz since 1980 using the 13.7 m radio telescope (Teraesranta et al., 1998). At present, there are 271 sources in the MRO monitoring program with data at 22 and 37 GHz, where the latter includes observations spanning up to 42 years. Data from this ongoing program have contributed to research on the long-term behavior, statistical properties, source populations, and radio variability models of AGNs. This monitoring has enabled studies of the radio to -ray connection in blazars, as well as radio versus optical behavior and multi-frequency variability models (e.g. Hovatta et al., 2007; Kankkunen et al., 2025). The Submillimeter Array (SMA; Gurwell et al., 2007) in Hawaii, United States, has been monitoring the quasars to determine the flux densities and variability at 230 and 345 GHz since the beginning of 2003 with 437 sources at present. The observations from this program were used in different long-term timing and spectral studies of AGNs (e.g. Villata et al., 2009; Fuhrmann et al., 2014; Jorstad and Marscher, 2016).
The above-mentioned radio monitoring programs were successful in providing data, complementing other multiband observations. The studies above show that a comprehensive investigation into radio and radio/-ray populations helps to address the enduring questions about AGNs physical attributes. These include the location, structure, and radiative characteristics of the -ray emission region. Additionally, such research addresses the collimation, composition, particle acceleration, and emission mechanisms in AGNs/blazar jets. Also, to detect the presence of shorter and longer time scales of quasi-periodic oscillations (QPOs) in the AGNs light curves, it is essential to have long-term adequate continuous data (e.g. King et al., 2013; Li et al., 2023).
Blazars have also been identified as candidate sources of PeV-energy neutrinos (IceCube Collaboration et al., 2018; Buson et al., 2022). To thoroughly explore blazars and the link between their radio emissions and neutrinos, multiband, high-cadence, long-term observations are essential (Kovalev et al., 2020; Plavin et al., 2020; Hovatta et al., 2021).
Motivated by this, the quasi-Simultaneous Multiwavelength Monitoring of gamma-ray-loud AGNs with the Nanshan 26-m radio telescope (SMMAN) program operated by Xinjiang Astronomical Observatory (XAO) in China started in late 2016 and remains active. The information obtained from the SMMAN program, along with Fermi-LAT observations, will enable to systematically determine the radio and its connection with other multiband observational characteristics of the gamma-ray-loud AGNs population. In the following, we discuss the SMMAN dataset in 4.8 and 23.6 GHz. The details of the source sample and the radio data analysis are described in Section II. In Section III, we present and discuss the results. Finally, we summarize and conclude the findings of our work in Section IV.
II Observation and data reduction
II.1 Facility
The SMMAN program is carried out with the NanShan 26-m Radio Telescope (NSRT), which is located on the Nanshan Mountain, with the geographic coordinates of 43°28’17”N, 87°10’41”E, and an altitude of 2080 meters, km southwest of Urumqi city, China. Apart from single-dish scientific observations, NSRT has been involved in a large number of Very Long Baseline Interferometry (VLBI) observations in the last years, as a formal member of the European VLBI Network (EVN). The NSRT has an on-axis, single-beam optical configuration and four receivers operating at L-band (1-2 GHz), S/X-band (2.3/8.6 GHz), C-band (4-8 GHz), and K-band (22-24 GHz). The receivers used in the SMMAN program operate in the C and K bands, centered at 4800 MHz and 23600 MHz, with terminal recording bandwidths of 500 MHz for both. The main technical specifications of the receivers used in the SMMAN program are listed in Table 1.
| Parameters | C-band | K-band | |
| Center frequency | [MHz] | 4800 | 23600 |
| Terminal bandwidth | [MHz] | 500 | 500 |
| System temperature | [K] | 26 | 45 |
| Reference signal | [K] | 1.7 | 2.5 |
| FWHP/Beam | [arcsec] | 546 | 105 |
| Aperture efficiency | [%] | 65 | 62 |
| Conversion Factor | [K/Jy] | 0.174 | 0.208 |
II.2 Source selection and monitoring schedule
Our goal is to monitor those bright variable blazars, the majority of which have been captured by the Fermi-LAT. Therefore, the initial source sample was established by filtering the Fermi-LAT 3FGL catalog (Acero et al., 2015) and cross-matching with the NRAO VLA Sky Survey (NVSS)222https://www.cv.nrao.edu/nvss/ by putting a limit on flux density at 1.4 GHz (, point-like structure, and source declination (J2000) .
Finally, our SMMAN sample contains 131 gamma-ray-loud AGNs, which include 113 blazars (FSRQs = 81 and BL Lacs = 32), nine active galaxies of uncertain type (bcu), seven radio galaxies (rdg), one compact steep spectrum quasar (css), and one steep spectrum radio quasar (ssrq), respectively.
The distribution of SMMAN sources in the sky and their redshift distribution is illustrated in Figure 1.
The redshift values were obtained from the NASA/IPAC Extragalactic Database (NED) and the SIMBAD astronomical database (Wenger et al., 2000). Redshift measurements are available for 124 of the 131 sources in our monitored sample, of which 121 are spectroscopic and three are photometric.
These sources were quasi-simultaneously monitored in the C and K bands using the 26-meter NSRT for the last approximately eight years (2016-2024), with a cadence of about one month.
II.3 Data Acquisition and Reduction
Flux density measurements with single-dish antennas can be performed through multiple modes, such as on-off, mapping, and cross-scan. Considering the monitoring efficiency and measurement accuracy, our SMMAN program has adopted the cross-scan mode. The cross-scan mode is achieved by performing multiple sub-scans of the target source in two orthogonal directions. Our observations were carried out with 4+4 sub-scans in azimuth and elevation over the target’s position, respectively. The data were acquired in this manner at both C-band (4.8 GHz) and K-band (23.6 GHz). For a single sub-scan, the scan width is set to 5 times the beam width, which is and arcmin for C-band and K-band, respectively. The total sampling time is set to 20 to 30 seconds for each sub-scan, while the sampling time for each data point is fixed at 64 ms during scanning.
The target samples are mainly point sources for the 26-meter NSRT at C and K bands. A sub-scan across the target is simply equal to a beam-wide Gaussian, which is the convolution of the source and the telescope beam. Given the structural perturbations induced by the telescope scanning motions, a Gaussian kernel with a full width at half maximum (FWHM) comparable to the beam size was adopted for scan data model fitting. During cross-scanning, the response attenuation in the orthogonal direction is determined by the Gaussian peak offset along both scan axes, thereby allowing for correcting flux density measurement reduction caused by pointing deviation. As an example, such a cross-scan mode is illustrated in Figure 2.
The data reduction and calibration process includes Gaussian fitting, pointing correction, gain correction, and absolute flux density calibration. Gaussian fitting is used to subtract the baseline of the system temperature drifting and extract the Gaussian-like signals from the sub-scans. Pointing correction is to calibrate the antenna pointing error with each other direction’s offset in the scans. After that, we subsequently conducted atmospheric opacity and antenna gain calibration to further correct the elevation gain effect. Note that the systematic error possibly arising from solar elongation dependence was also accounted for during the pointing correction step. Finally, the absolute flux density was tuned using the calibration factors generated by three primary calibrators, 3C 48, NGC 7027, and 3C 286 (Baars et al., 1977; Ott et al., 1994; Zijlstra et al., 2008; Angelakis et al., 2015). The observed average flux density for these calibrator sources during the SMMAN program is presented in Table 2.
| Source | |||
| (GHz) | 3C 48 | 3C 286 | NGC 7027 |
| 4.8 | 5.63 (0.11) | 7.51 (0.16) | 5.37 (0.11) |
| 23.6 | 1.26 (0.19) | 2.43 (0.22) | 5.40 (0.39) |
All the data reduction and calibration processes above are conducted with the NSRT radio continuum data processing software Data Analysis Application for Radio Continuum (DAARC)333https://github.com/junliu/daarc, which is developed by our team in Python language.
II.4 -ray data
We used the monthly binned -ray flux and spectral index data from the Fermi-LAT Light Curve Repository444https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/ (LCR; Abdollahi et al. 2023) in the energy range of GeV, covering the same time period as the radio observations (2016 November 1 to 2024 November 30) presented in this study. For a detailed description of the unbinned maximum-likelihood data analysis, we refer to Abdollahi et al. (2020). In this work, we used the light curve obtained by allowing the spectral slope to vary freely during the fitting process and the fit results of only the convergent likelihood analysis.
The significance of a source detection is evaluated using the test statistic , where is the likelihood of the null model (no source) and is the likelihood of the model that includes the source. In the monthly binned light curve, we only consider those bins with detection significance of (), and the upper limit values (for non-detection bins) were not included. It is important to note that performing a full Fermi-LAT data analysis for 131 SMMAN sources is computationally demanding and time-consuming. Therefore, we used the available light curves from the LCR. These light curves of various -ray sources from this repository have been used in previous research studies (e.g. Di Gesu et al., 2023; Wang et al., 2023).
III Results and Discussion
The basic information about the sources that were monitored in this work is listed in Table 3 of Appendix A.
Note that in our analysis and discussion part, we have considered the different AGN classes from Table 3 of the Appendix A as FSRQ (FSRQ+fsrq), BL Lac (BLL+bll), BCU (bcu), and RDG (RDG+rdg+css+ssrq). Here, the AGNs class is defined according to the Fermi Large Area Telescope Fourth Source Catalog Data Release 4 (4FGL-DR4; Ballet et al. 2023), with designations in capital letters representing confirmed identifications, and those in lower-case letters denoting associations.
The number of radio measurements for the SMMAN sources between 1 November 2016 and 30 November 2024, as presented in this work, is shown in Figure 3.
Most sources have more observations at 23.6 GHz than at 4.8 GHz, and some sources have exceptionally large datasets with more than 400 measurements.
All histogram figures presented in this work use binning based on Knuth’s optimal, data-driven method (Knuth, 2006). The overlap of the SMMAN sources with other historical surveys or ongoing monitoring programs is displayed in Figure 4.
From this plot, one can easily identify the number of sources having counter monitoring data for the SMMAN sample, which is useful for further studies that could cover more multiband radio observations.
The sources that overlap among the SMMAN targets and UMRAO, F-GAMMA, RATAN-600, OVRO, ROBIN, MRO, and SMA monitoring programs are shown in Table 4 of Appendix A.
To depict the importance of the SMMAN monitoring program, a long-term multiband light curve of the BL Lac object S4 0954+65 using the data from the SMMAN, RATAN-600555https://www.sao.ru/blcat/(Mingaliev et al., 2014; Sotnikova et al., 2022), F-GAMMA (Angelakis et al., 2019a, b) and ROBIN (Marchili et al., 2025a, b) monitoring program is shown in Figure 5. The plot clearly shows that the SMMAN program has covered a greater number of observations after 2016 November compared to the other surveys (publicly available), which is important due to the fact that adequate continuous data are essential for long-term temporal and spectral studies, and the detection of QPOs. In addition, a part of the data from the SMMAN program has been used in systematic studies of a few individual blazars in earlier research works (e.g. Kun et al., 2022; Krishna Mohana et al., 2024; Kun et al., 2024; Mohana A et al., 2025), which also showcase the significance of the SMMAN program along with other multiwavelength observations.
III.1 Variability and Other Physical Parameter Estimations
To quantify the variability of sources, at each frequency band, the variability index () is calculated using the formula adopted from Aller et al. (1992), defined as
where and are the maximum and minimum values of the flux density over all epochs of observations, while and are their root-mean-square errors. The uncertainty of the variability index was calculated as
The variability index is based solely on the two extreme flux density values. Therefore, it is necessary to characterize the variability of the light curve as a whole, incorporating all available flux density measurements. Hence, to identify the presence of true intrinsic variability in the light curve, we have used the fractional root mean square variability parameter (), since it considers the uncertainties in flux density measurements. Here we use the relation given in Vaughan et al. (2003), which is defined as:
where, = S2 – , is called excess variance, S2 is the sample variance, is the mean square uncertainties of each observation and is the sample mean.
The radio spectral index is calculated in two frequency intervals (4.8 and 23.6 GHz). The radio spectral index defined from the power law , where is the flux density at frequency , and is a spectral slope. It is calculated using the formula:
where and are the flux densities at frequencies and , respectively. The uncertainty in the spectral index was determined using Gaussian error propagation as:
where the factor of arises from the derivation of the spectral index as defined in Equation 5.
The radio luminosity was calculated at 4.8 and 23.6 GHz following the equation:
where is the measured flux density at 4.8 or 23.6 GHz, is the redshift, is the average spectral index in the interval 4.8 and 23.6 GHz, and is the luminosity distance. For calculation of luminosity distances, we used the CDM cosmology with km s-1 Mpc-1, , and (Planck Collaboration et al., 2016).
III.2 Characteristics of Radio Variability, Spectral index and Luminosity: Relation to AGNs Class
The estimated variability index values for different sources at 4.8 and 23.6 GHz are presented in
Table 5 of Appendix A. The histograms of
at two frequencies are shown in Figure 6.
At 23.6 GHz of the sources exhibit variability above the level (), whereas at lower frequency 4.8 GHz sources show significant variability. The range in the variability for FSRQs and BL Lacs at 4.8 GHz is comparable (varying from up to ) with the mean value for FSRQs at and for BL Lacs at . The BCU and RDG classes vary from up to maximum values of and , respectively. Similarly, at 23.6 GHz, the range of variability in FSRQs and BL Lacs is comparable (varying from up to ) with the mean value for FSRQs at and for BL Lacs at . The BCU and RDG classes vary from up to maximum values of and , respectively.
The calculated fractional variability values for different
SMMAN sources in the C and K bands are given in Table 5 of the Appendix A. In Figure 7, we show fractional variability histograms at 4.8 and 23.6 GHz for different classes of AGNs in our sample. It is observed that at 4.8 GHz, of the sources show fractional variability
. The value of the FSRQs found vary from to a maximum of with a mean value of , BL Lacs show a range between and a maximum value of with a mean value of .
A similar trend of variation is observed in active galaxies of uncertain type with a maximum value up to and in radio galaxies with maximum variation up to . In contrast to lower frequency behavior, of the sources exhibit fractional variability . Here, the different classes show variation of to maximum values of , , , and in FSRQs, BL Lacs, BCUs, and RDGs, respectively. At 23.6 GHz, the mean value of for FSRQs and BL Lacs is found to be and , respectively.
It is noticed that from both the estimation of variability index and fractional variability values at 4.8 and 23.6 GHz, a trend for sources to exhibit a higher degree of variability at 23.6 GHz than at 4.8 GHz, with a mean value for BL Lacs higher than the FSRQs. These results are consistent with the pattern reported by Aller et al. (1992) and Sotnikova et al. (2022) for a sample comprising 62 extragalactic objects (at 4.8 and 14.5 GHz) and 1700 blazars (in frequency range 1.2-22.3 GHz), respectively.
The average spectral index with associated error and the range in spectral index (minimum and maximum values) over the entire observing period are presented in Table 5 of the Appendix A. It is worth noting that when estimating the spectral index, monthly averaged flux densities were used to ensure temporal alignment between the two frequencies.
In Figure 8, the left panel shows the distribution of average spectral indices, while the right panel displays the range of spectral indices versus their average values for the various AGNs classes in the SMMAN sample. BL Lac objects tend to exhibit comparatively flatter spectra than the other classes, whereas FSRQs, BCUs, and RDGs show a broader distribution, spanning from flat to steep.
The mean spectral indices for FSRQs is with values ranging between and . For BL Lacs, the spectral index ranges from to , with a mean of . Radio galaxies generally display steep spectra (ranging between to ), and the BCUs are also skewed toward steeper spectral indices (varying from to ). Also, the BL Lacs and FSRQs were found to show a higher degree of variability trend at both frequency bands compared to the other classes.
Figure 9 shows variability indices versus average spectral index in the range of 4.8-23.6 GHz. At both 4.8 and 23.6 GHz, of the sources with show variability greater than level. The majority of these sources belong to the FSRQ and BL Lac classes. Also, sources with exhibit a low degree of variability at both frequency bands. Interestingly, the radio galaxy 3C 111 with shows variability above at both 4.8 and 23.6 GHz with a maximum value of at 23.6 GHz. Such large variability could be due to the major outburst, which is usually associated with the release of radio knots from the core (Kadler et al., 2008; Grossberger et al., 2012). Similar features are noticed in the case of fractional variability versus the average spectral index plot (Figure 10).
The radio luminosity was calculated using the
Equation 7 for different classes of AGNs in our sample, the distribution is presented in Figure 11, and the values are given in the
Table 5 of the Appendix A. It should be noted that, among the redshifts used in our calculations, three out of the 124 measurements () are photometric, while the remaining measurements are spectroscopic.
The distribution ranges from to W Hz-1 at 4.8 GHz, and from to W Hz-1 at 23.6 GHz. In both frequency bands, FSRQs exhibit higher radio luminosity than BL Lacs and other classes. Specifically, at 4.8 GHz, FSRQs have radio luminosities from to W Hz-1 (mean , W Hz-1), while BL Lacs range from to W Hz-1 (mean , W Hz-1). At 23.6 GHz, the BL Lacs span from to W Hz-1 (mean , W Hz-1) and FSRQs from to W Hz-1 (mean , W Hz-1). In both bands, BCUs have radio luminosities intermediate between BL Lacs and FSRQs. RDGs exhibit lower luminosities, typically below the mean value of FSRQs.
III.3 Gamma-ray and Radio Properties of the SMMAN Sample
To investigate the connection between radio and -ray emissions, we use Fermi-LAT data in the energy range of GeV covering the same time period (2016 November 1 to 2024 November 30) as that of the radio observation presented in this work. The details of the -ray data are provided in Section II. The LCR provides light curves for all sources in the 4FGL-DR2 catalog with a variability index greater than 21.67 (Abdollahi et al., 2023). This index serves as a proxy for the average fractional variability , where is measured on timescales of one year. Consequently, among the 131 SMMAN sources, we obtained monthly binned light curves and spectral indices for 105. From the light curves, we calculated the average flux and spectral indices; the values are presented in Table 5 of the Appendix A. Further, following the procedures described in Böck et al. (2016) with the assumption that sources have a power-law photon flux spectrum (Equation 8), the -ray energy flux and luminosities for the 105 sources were calculated as follows.
where is the estimated photon flux in the energy band from to ( is the reference energy used to make the expression dimensionless while applying a non-integer exponent), for photon index . Then the energy flux in that band is given by
for .
Because the sources in the sample are distributed across various redshifts, the measured luminosities were corrected using the K-correction as described by Ghisellini et al. (2009), i.e.,
where is the redshift of the source, the energy
flux, and the photon spectral index in the -ray band.
The luminosity distance, , was calculated assuming a
flat universe with the cosmological parameters same as those used in the estimation of the radio luminosity.
The -ray flux distribution for the SMMAN sample is shown in the top panel of Figure 12, while the bottom panel presents the corresponding distribution of spectral indices. On average, BL Lacs tend to exhibit harder spectra than FSRQs, consistent with earlier results based on the 3FGL sources (see Figure 7 of Ackermann et al. 2015). We performed the two sample
Kolmogorov–Smirnov (KS) test to check whether the BL Lac objects have the same distribution as those of FSRQs. The test shows a strong statistical result with a -value of .
This result is in agreement with earlier studies, which have shown a relationship between the -ray spectral index and the peak frequency of the synchrotron component in the spectral energy distribution.
It is to be noted that blazars were also classified into three subclass based on the synchrotron peak frequency (): low-synchrotron-peaked (LSP; ), intermediate-synchrotron-peaked (ISP; ), and high-synchrotron-peaked (HSP; ) blazars (Abdo et al., 2010a; Fan et al., 2016; Yang et al., 2022).
As the synchrotron peak frequency drops, their -ray spectra become softer (see, e.g., Figure 17 of Ackermann et al. (2011) and Figure 10 of Ackermann et al. (2015)).
The observed -ray luminosities, as calculated from Equation 10 and given in Table 5 of the Appendix A, exhibit a clear dependence on the classification of the source (top panel of Figure 13).
The luminosities of radio galaxies are low, those of BL Lac objects are in between, while FSRQs exhibit high luminosities. The bottom panel of Figure 13 shows the relation between the -ray luminosity and the spectral index. A trend of the spectral index becoming harder with decreasing luminosity is noticed. This relation for all Fermi-LAT detected
AGNs in 3FGL is presented in Figure 14 of Ackermann et al. (2015) and in the Figure 10 of Ajello et al. (2020) for the fourth Fermi-Large Area Telescope Source Catalog (4FGL; Abdollahi et al. 2020).
This pattern has been widely addressed in the perspective of “blazar sequence” in the earlier studies (Ackermann et al. 2015, and references therein).
Figure 14 shows the radio luminosity at 23.6 GHz as a function of redshift (top panel) and the -ray luminosity as a function of redshift (bottom panel) for the sample of SMMAN sources having radio and -ray luminosities. Clearly, a selection effect is evident: sources with higher luminosities are preferentially detected at higher redshifts in both the radio and -ray bands. FSRQs exhibit higher radio and -ray luminosities and are predominantly found at higher redshifts, whereas BL Lacs tend to have lower radio and -ray luminosities and are generally located at lower redshifts. Radio galaxies span a broader range, extending from the lowest luminosities and redshifts to intermediate values between the FSRQ and BL Lac regimes. Such trend is also noticed for the radio luminosity at 4.8 GHz as a function of redshift.
Figure 15 shows the -ray flux and radio flux density relation (top panel for 4.8 GHz and bottom panel for 23.6 GHz). The Kendall’s rank correlation coefficient is 0.29 with a -value of and 0.37 with a -value of for the cases of 4.8 and 23.6 GHz, respectively, thus confirming the correlation between these quantities. The -ray luminosities were calculated as explained in Section II. The
relation between -ray and radio luminosity is presented in Figure 16.
A linear fit in log-log space gives (Pearson correlation coefficient = 0.68) and (Pearson correlation coefficient = 0.86) for radio luminosities of 4.8 and 23.6 GHz, respectively. Such a correlation could be induced by observational effects, which depend on the source distance (e.g., Böck et al., 2016). This is further supported by Figure 14, which indicates that the observed correlation in Figure 16 is likely largely driven by the fact that our sample is both radio flux-limited and -ray flux-limited. Consequently, a strong correlation is expected as a result of Malmquist bias.
Following the definition of -ray loudness in Böck et al. (2016)
as the ratio of the total -ray flux to the radio flux density
(), the distribution of -ray loudness is shown in Figure 17 for both C and K band data.
Taking into account the case of both frequency bands, the -ray loudness value ranges from to . It is noted that both FSRQ and BL Lac classes have a higher loudness factor compared to radio galaxies, and the value for BCUs falls between them. The mean value of -ray loudness for FSRQs and BL Lacs is at and , respectively, in the C band case. For the K band, the mean values are and for FSRQs and BL Lacs, respectively.
The distribution of the -ray loudness shows a similar dependence on the source class as the -ray spectral index distribution.
Both BL Lacs and FSRQs have a wider distribution range; however, BL Lacs appear to have slightly higher -ray loudness factor compared to FSRQs. This relationship is consistent with shifts in the peak frequencies of the SEDs. When the synchrotron peak moves to higher frequencies, the radio-band flux density decreases. In contrast, a shift in the high-energy SED peak to higher frequencies increases the observed -ray flux and results in a more intense spectrum. This interpretation is further supported by the observed weak anticorrelation (Kendall rank correlation = -0.17 at 4.8 GHz and -0.19 at 23 GHz) between the -ray spectral index and the -ray loudness as shown in Figure 18. The weak anticorrelation observed here may be due to the consideration of the average -ray spectral index value calculated from the -ray monthly light curve taken from the Fermi LCR. This pattern is in agreement with the study reported by Böck et al. (2016) with a strong anticorrelation (see Figure 9).
IV Summary and Conclusions
We present a first release of a dataset of 131 gamma-ray-loud AGNs monitored using the Nanshan 26-m Radio Telescope (NSRT) during the November 2016 to November 2024 period in the frequency band of 4.8 and 23.6 GHz. We summarize and conclude our findings as follows.
1.
The main objective of the SMMAN program was primarily to offer essential multi-frequency radio monitoring at the lower frequency bands that complements the -ray observations conducted by Fermi-LAT along with the other multiband monitoring program and to investigate all pertinent aspects of radio physics of gamma-ray-loud AGNs.
2.
The light curves presented here are unbinned, and each data point was subjected to proper post-measurement corrections, including Gaussian fitting, pointing calibration, atmospheric opacity, gain calibration, and absolute flux density calibration. The complete dataset has undergone a final quality check. The light curves of 131 SMMAN sources and three calibrator sources are presented in Figure 19 of Appendix B.
3.
From the variability analysis, we find a trend in which sources show a higher degree of variability at 23.6 GHz than at 4.8 GHz, with BL Lacs exhibiting higher mean variability levels than FSRQs. BCUs and radio galaxies show lower variability compared to BL Lacs and FSRQs. In both frequency bands, FSRQs show higher radio luminosities than BL Lacs and other source classes. BCUs display radio luminosities intermediate between those of BL Lacs and FSRQs, while RDGs have lower luminosities, generally below the mean luminosity of FSRQs.
4.
The radio spectral index indicates that BL Lac objects generally have flatter spectra, whereas FSRQs, BCUs, and RDGs span a wider range from flat to steep. At 4.8 and 23.6 GHz, about 90% of sources with display variability above 10%, predominantly among FSRQs and BL Lacs. Sources with , however, show little variability at either frequency.
5.
The -ray loudness factor was estimated for different sources, and it was found that FSRQs and BL Lacs are louder than radio galaxies, whereas BCUs lie between these two populations. The distribution of -ray loudness exhibits a dependence on source class similar to that of the -ray spectral index distribution. In addition, a weak anticorrelation is observed between the -ray spectral index and the -ray loudness factor.
6.
The data release of long-term light curve from this monitoring program in the lower frequency bands provides a unique coverage of the radio part of the SED and is a good addition to the other historical and ongoing programs covering the frequency range from low to high frequency bands.
The users using the SMMAN radio data in their publications are kindly requested to acknowledge the source of the data by citing this paper. To use the data and for details, please contact Lang Cui (cuilang@xao.ac.cn).
Acknowledgements
We are thankful to Sándor Frey, Timur Mufakharov, and Junhui Fan for carefully reading the manuscript and providing insightful comments. We also thank the anonymous reviewer for valuable comments, which helped us to improve the manuscript. This work was supported by the Tianshan Talent Training Program (grant No. 2023TSYCCX0099) and the National Key R&D Program of China (grant Nos. 2024YFA1611500 and 2022SKA0120102). This work was also partly supported by the Urumqi Nanshan Astronomy and Deep Space Exploration Observation and Research Station of Xinjiang (XJYWZ2303) and the Central Guidance for Local Science and Technology Development Fund (grant No. ZYYD2026JD01). The Nanshan 26-m Radio Telescope (NSRT) is operated by the Xinjiang Astronomical Observatory (XAO), CAS. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has used the RATAN-600 data, provided by the Special Astrophysical Observatory of the Russian Academy of Sciences (SAO RAS).
Appendix A Source sample
Table 3 lists all the sources included in the current data release. In Table 4, the information on whether the data for SMMAN sources is available in the other historical or the ongoing monitoring programs, such as UMRAO, F-GAMMA, RATAN-600, OVRO, ROBIN, MRO, and SMA, is provided with the corresponding name in that monitoring program.
| Name 3FGL | Assoc. Name | Coordinates (J2000) | AGN | z | 4.8 GHz | 23.6 GHz | |||
| RA | Dec | Class | |||||||
| (Jy) | (Jy) | ||||||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) |
| J0012.4+7040 | TXS 0008+704 | 00 11 31.90 | +70 45 31.62 | bcu | … | 163 | 0.75 (0.25) | 68 | 0.88 (0.30) |
| J0014.6+6119 | 4C +60.01 | 00 14 48.79 | +61 17 43.54 | bcu | … | 226 | 1.97 (0.07) | 52 | 0.83 (0.10) |
| J0059.6+0003 | PKS 0056-00 | 00 59 05.51 | +00 06 51.62 | fsrq | 0.719 | 100 | 1.39 (0.11) | … | … |
| J0102.8+5825 | TXS 0059+581 | 01 02 45.76 | +58 24 11.14 | fsrq | 0.644 | 227 | 2.43 (0.68) | 227 | 3.59 (1.21) |
| J0108.7+0134 | 4C +01.02 | 01 08 38.77 | +01 35 00.32 | fsrq | 2.099 | 146 | 3.88 (1.27) | 157 | 5.66 (2.44) |
| J0110.2+6806 | 4C +67.04 | 01 10 12.87 | +68 05 41.22 | bll | 0.290 | 173 | 0.89 (0.05) | … | … |
| J0130.8+1441 | 4C +14.06 | 01 29 55.35 | +14 46 47.84 | fsrq | 1.630 | 39 | 0.57 (0.05) | … | … |
| J0137.0+4752 | OC 457 | 01 36 58.59 | +47 51 29.10 | fsrq | 0.859 | 201 | 1.91 (0.23) | 171 | 2.63 (0.51) |
| J0205.0+1510 | 4C +15.05 | 02 04 50.41 | +15 14 11.04 | bcu | 0.834 | 157 | 2.50 (0.42) | 93 | 1.60 (0.69) |
| J0217.5+7349 | S5 0212+73 | 02 17 30.81 | +73 49 32.62 | fsrq | 2.346 | 211 | 3.42 (0.12) | 226 | 2.62 (0.56) |
| J0221.1+3556 | B0218+357 | 02 21 05.47 | +35 56 13.72 | FSRQ | 0.944 | 124 | 1.03 (0.08) | 15 | 0.68 (0.08) |
| J0222.6+4301 | 3C 66A | 02 22 39.61 | +43 02 07.80 | BLL | 0.370 | … | … | 5 | 0.85 (0.20) |
| J0237.9+2848 | 4C +28.07 | 02 37 52.41 | +28 48 08.99 | FSRQ | 1.213 | 175 | 2.43 (0.44) | 140 | 2.19 (0.46) |
| J0238.6+1636 | AO 0235+164 | 02 38 38.93 | +16 36 59.28 | BLL | 0.940 | 284 | 1.49 (0.48) | 213 | 1.74 (0.45) |
| J0242.3+1059 | OD 166 | 02 42 29.17 | +11 01 00.73 | fsrq | 2.680 | 52 | 0.84 (0.04) | 29 | 0.74 (0.18) |
| J0303.6+4716 | 4C +47.08 | 03 03 35.24 | +47 16 16.27 | bll | 0.475 | 184 | 1.85 (0.58) | 105 | 1.73 (0.64) |
| J0308.6+0408 | NGC 1218 | 03 08 26.22 | +04 06 39.30 | rdg | 0.028 | 144 | 3.61 (0.10) | 83 | 1.16 (0.13) |
| J0319.8+4130 | NGC 1275 | 03 19 48.16 | +41 30 42.11 | RDG | 0.018 | 140 | 32.46 (5.45) | 145 | 30.08 (5.15) |
| J0336.5+3210 | NRAO 140 | 03 36 30.11 | +32 18 29.34 | fsrq | 1.265 | 156 | 1.91 (0.22) | 117 | 1.42 (0.23) |
| J0354.1+4643 | B3 0350+465 | 03 54 30.01 | +46 43 18.75 | bcu | … | 56 | 0.72 (0.13) | 16 | 0.57 (0.15) |
| J0358.8+6002 | TXS 0354+599 | 03 59 02.64 | +60 05 22.07 | fsrq | 0.455 | 181 | 1.12 (0.09) | 59 | 0.84 (0.14) |
| J0418.5+3813c | 3C 111 | 04 18 21.28 | +38 01 35.80 | rdg | 0.048 | 275 | 6.86 (0.39) | 271 | 2.41 (0.81) |
| J0423.8+4150 | 4C +41.11 | 04 23 56.01 | +41 50 02.71 | bll | 129 | 1.58 (0.24) | 83 | 1.18 (0.22) | |
| J0510.0+1802 | PKS 0507+17 | 05 10 02.37 | +18 00 41.58 | fsrq | 0.416 | 117 | 1.02 (0.12) | 117 | 2.13 (0.42) |
| J0512.9+4038 | B3 0509+406 | 05 12 52.54 | +40 41 43.62 | bcu | … | 140 | 1.07 (0.21) | 80 | 1.05 (0.33) |
| J0517.4+4540 | 4C +45.08 | 05 17 28.90 | +45 37 04.86 | FSRQ | 0.839 | 101 | 0.79 (0.08) | … | … |
| J0519.3+2746 | 4C +27.15 | 05 19 33.02 | +27 44 04.27 | bcu | 106 | 0.72 (0.04) | … | … | |
| J0521.7+2113 | TXS 0518+211 | 05 21 45.96 | +21 12 51.45 | bll | 0.108 | 27 | 0.50 (0.09) | … | … |
| J0532.7+0732 | OG 050 | 05 32 39.00 | +07 32 43.34 | FSRQ | 1.254 | 125 | 1.67 (0.20) | 95 | 1.72 (0.34) |
| J0533.2+4822 | TXS 0529+483 | 05 33 15.86 | +48 22 52.81 | fsrq | 1.160 | 105 | 1.13 (0.31) | 68 | 1.27 (0.44) |
| J0603.8+2155 | 4C +22.12 | 06 03 51.56 | +21 59 37.70 | bcu | … | 113 | 1.29 (0.05) | 32 | 0.60 (0.15) |
| J0638.6+7324 | S5 0633+73 | 06 39 21.96 | +73 24 58.04 | fsrq | 1.850 | 161 | 0.99 (0.13) | 121 | 0.96 (0.27) |
| J0710.5+4732 | S4 0707+47 | 07 10 46.10 | +47 32 11.14 | bll | 1.292 | 74 | 0.61 (0.06) | … | … |
| J0721.9+7120 | S5 0716+71 | 07 21 53.45 | +71 20 36.36 | BLL | 0.230 | 648 | 1.25 (0.27) | 626 | 1.56 (0.80) |
| J0750.6+1232 | OI 280 | 07 50 52.05 | +12 31 04.83 | fsrq | 0.889 | 127 | 1.91 (0.32) | 95 | 1.95 (0.41) |
| J0757.0+0956 | PKS 0754+100 | 07 57 06.64 | +09 56 34.85 | bll | 0.266 | 126 | 1.21 (0.22) | 84 | 1.20 (0.26) |
| J0758.7+3747 | NGC 2484 | 07 58 28.11 | +37 47 11.81 | rdg | 0.041 | 103 | 1.22 (0.04) | … | … |
| J0805.4+6144 | TXS 0800+618 | 08 05 18.18 | +61 44 23.70 | fsrq | 3.018 | 66 | 0.66 (0.09) | … | … |
| J0809.5+4045 | S4 0805+41 | 08 08 56.65 | +40 52 44.89 | fsrq | 1.418 | 118 | 1.08 (0.14) | 3 | 1.48 (0.18) |
| J0818.2+4223 | S4 0814+42 | 08 18 16.00 | +42 22 45.42 | bll | 0.530 | 120 | 1.30 (0.28) | … | … |
| J0824.9+5551 | OJ 535 | 08 24 47.24 | +55 52 42.67 | fsrq | 1.419 | 166 | 1.20 (0.13) | … | … |
| J0824.9+3916 | 4C +39.23 | 08 24 55.48 | +39 16 41.91 | fsrq | 1.216 | 93 | 1.24 (0.08) | 31 | 0.90 (0.24) |
| J0830.7+2408 | OJ 248 | 08 30 52.09 | +24 10 59.82 | FSRQ | 0.939 | 99 | 0.95 (0.11) | 35 | 0.95 (0.22) |
; the number of observations in 4.8 and 23.6 GHz for each source (Col. 7 and 9), and the average flux density calculated over these observations (Col. 8 and 10).
| Name 3FGL | Assoc. Name | Coordinates (J2000) | AGN | z | 4.8 GHz | 23.6 GHz | |||
| RA | Dec | Class | |||||||
| (Jy) | (Jy) | ||||||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) |
| J0832.6+4914 | OJ 448 | 08 32 23.22 | +49 13 21.04 | bll | 0.548 | 43 | 0.56 (0.10) | … | … |
| J0840.8+1315 | 3C 207 | 08 40 47.59 | +13 12 23.56 | ssrq | 0.680 | 123 | 1.83 (0.17) | 89 | 1.62 (0.43) |
| J0841.4+7053 | 0836+710 | 08 41 24.37 | +70 53 42.17 | FSRQ | 2.213 | 889 | 2.88 (0.14) | 776 | 2.01 (0.46) |
| J0854.8+2006 | OJ 287 | 08 54 48.87 | +20 06 30.64 | BLL | 0.306 | 138 | 4.40 (0.74) | 126 | 6.40 (1.12) |
| J0903.1+4649 | S4 0859+47 | 09 03 03.99 | +46 51 04.14 | fsrq | 1.465 | 139 | 1.39 (0.08) | 77 | 0.96 (0.23) |
| J0904.3+4240 | S4 0900+42 | 09 04 15.63 | +42 38 04.76 | fsrq | 1.342 | 61 | 0.85 (0.08) | … | … |
| J0909.1+0121 | PKS 0906+01 | 09 09 10.09 | +01 21 35.62 | fsrq | 1.023 | 86 | 1.60 (0.21) | 77 | 1.58 (0.48) |
| J0916.3+3857 | S4 0913+39 | 09 16 48.90 | +38 54 28.15 | fsrq | 1.268 | 58 | 0.80 (0.06) | … | … |
| J0920.9+4442 | S4 0917+44 | 09 20 58.46 | +44 41 53.98 | fsrq | 2.188 | 123 | 1.34 (0.13) | 132 | 2.04 (0.51) |
| J0921.8+6215 | OK 630 | 09 21 36.23 | +62 15 52.18 | fsrq | 1.458 | 172 | 1.10 (0.13) | 77 | 1.14 (0.28) |
| J0948.6+4041 | 4C +40.24 | 09 48 55.34 | +40 39 44.58 | fsrq | 1.250 | 112 | 1.22 (0.11) | … | … |
| J0957.6+5523 | 4C +55.17 | 09 57 38.18 | +55 22 57.77 | fsrq | 0.901 | 159 | 1.83 (0.05) | 55 | 0.96 (0.20) |
| J0957.4+4728 | OK 492 | 09 58 19.67 | +47 25 07.84 | fsrq | 1.884 | 115 | 1.16 (0.25) | 38 | 1.03 (0.19) |
| J0958.6+6534 | S4 0954+65 | 09 58 47.25 | +65 33 54.82 | BLL | 0.369 | 542 | 1.04 (0.32) | 565 | 1.72 (0.90) |
| J1023.1+3952 | 4C +40.25 | 10 23 11.57 | +39 48 15.38 | fsrq | 1.254 | 86 | 0.72 (0.07) | … | … |
| J1033.2+4116 | S4 1030+41 | 10 33 03.71 | +41 16 06.23 | fsrq | 1.117 | 115 | 1.30 (0.45) | 57 | 1.05 (0.26) |
| J1033.8+6051 | S4 1030+61 | 10 33 51.43 | +60 51 07.33 | FSRQ | 1.408 | 73 | 0.67 (0.13) | … | … |
| J1043.1+2407 | B2 1040+24A | 10 43 09.04 | +24 08 35.41 | fsrq | 0.562 | 84 | 0.79 (0.17) | 74 | 0.92 (0.32) |
| J1044.4+8058 | S5 1039+81 | 10 44 23.06 | +80 54 39.44 | fsrq | 0.168 | 175 | 1.20 (0.16) | 141 | 1.05 (0.22) |
| J1048.4+7144 | S5 1044+71 | 10 48 27.62 | +71 43 35.94 | FSRQ | 1.140 | 206 | 2.01 (0.71) | 235 | 2.11 (0.64) |
| J1058.5+0133 | 4C +01.28 | 10 58 29.61 | +01 33 58.82 | BLL | 0.894 | 106 | 4.75 (0.59) | 93 | 6.13 (0.99) |
| J1128.0+5921 | TXS 1125+596 | 11 28 13.34 | +59 25 14.80 | fsrq | 1.799 | 159 | 1.08 (0.13) | 109 | 0.93 (0.24) |
| J1131.4+3819 | B2 1128+38 | 11 30 53.28 | +38 15 18.55 | fsrq | 1.740 | 120 | 1.75 (0.31) | 36 | 1.91 (0.47) |
| J1132.8+1015 | 4C +10.33 | 11 32 59.49 | +10 23 42.26 | FSRQ | 0.540 | 16 | 0.45 (0.04) | … | … |
| J1145.1+1935 | 3C 264 | 11 45 05.01 | +19 36 22.74 | rdg | 0.022 | 121 | 2.04 (0.06) | … | … |
| J1146.8+3958 | S4 1144+40 | 11 46 58.30 | +39 58 34.30 | fsrq | 1.088 | 121 | 1.51 (0.26) | 12 | 2.49 (0.59) |
| J1153.4+4932 | OM 484 | 11 53 24.47 | +49 31 08.83 | FSRQ | 0.334 | 100 | 1.32 (0.18) | 71 | 1.08 (0.34) |
| J1153.4+4033 | B3 1151+408 | 11 53 54.66 | +40 36 52.62 | fsrq | 0.925 | 225 | 1.02 (0.10) | … | … |
| J1159.5+2914 | Ton 599 | 11 59 31.83 | +29 14 43.83 | fsrq | 0.725 | 142 | 2.55 (1.09) | 139 | 4.07 (1.69) |
| J1213.7+1306 | 4C +13.46 | 12 13 32.16 | +13 07 20.56 | fsrq | 1.140 | 42 | 0.74 (0.06) | … | … |
| J1224.9+2122 | 4C +21.35 | 12 24 54.46 | +21 22 46.39 | FSRQ | 0.434 | 122 | 2.15 (0.49) | 81 | 1.34 (0.46) |
| J1229.1+0202 | 3C 273 | 12 29 06.70 | +02 03 08.60 | FSRQ | 0.158 | 282 | 29.12 (0.85) | 421 | 17.07 (2.17) |
| J1230.9+1224 | M 87 | 12 30 49.42 | +12 23 28.04 | rdg | 0.004 | 122 | 68.46 (1.35) | 118 | 17.49 (1.32) |
| J1239.4+0727 | PKS 1236+077 | 12 39 24.59 | +07 30 17.19 | bll | 0.400 | 62 | 0.74 (0.11) | … | … |
| J1256.1-0547 | 3C 279 | 12 56 11.17 | -05 47 21.52 | FSRQ | 0.536 | 178 | 13.03 (0.89) | 335 | 20.32 (5.12) |
| J1258.1+3233 | ON 393 | 12 57 57.23 | +32 29 29.33 | fsrq | 0.806 | 42 | 0.66 (0.15) | … | … |
| J1302.6+5748 | TXS 1300+580 | 13 02 52.47 | +57 48 37.61 | bll | 0.950 | 43 | 0.62 (0.14) | … | … |
| J1309.5+1154 | 4C +12.46 | 13 09 33.93 | +11 54 24.55 | BLL | 1.453 | 41 | 0.93 (0.07) | 22 | 0.87 (0.12) |
| J1310.6+3222 | OP 313 | 13 10 28.66 | +32 20 43.78 | fsrq | 1.678 | 133 | 1.72 (0.60) | 64 | 3.30 (0.69) |
| J1351.1+0030 | PKS 1348+007 | 13 51 04.43 | +00 31 19.40 | fsrq | 2.084 | 4 | 0.54 (0.22) | … | … |
| J1419.9+5425 | OQ 530 | 14 19 46.60 | +54 23 14.79 | bll | 0.152 | 124 | 1.25 (0.48) | 71 | 1.06 (0.42) |
| J1442.6+5156 | 3C 303 | 14 43 02.78 | +52 01 37.27 | rdg | 0.141 | 113 | 1.16 (0.05) | 36 | 0.85 (0.19) |
| J1500.6+4750 | TXS 1459+480 | 15 00 48.65 | +47 51 15.54 | bll | 1.059 | 33 | 0.47 (0.09) | … | … |
| J1540.8+1449 | 4C +14.60 | 15 40 49.49 | +14 47 45.88 | bll | 0.606 | 109 | 1.24 (0.23) | … | … |
| J1549.4+0237 | PKS 1546+027 | 15 49 29.44 | +02 37 01.16 | fsrq | 0.414 | 106 | 2.86 (0.61) | … | … |
| Name 3FGL | Assoc. Name | Coordinates (J2000) | AGN | z | 4.8 GHz | 23.6 GHz | |||
| RA | Dec | Class | |||||||
| (Jy) | (Jy) | ||||||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) |
| J1550.5+0526 | 4C +05.64 | 15 50 35.27 | +05 27 10.45 | fsrq | 1.420 | 98 | 2.30 (0.08) | … | … |
| J1604.6+5714 | GB6 J1604+5714 | 16 04 37.35 | +57 14 36.66 | fsrq | 0.720 | … | … | 40 | 0.80 (0.25) |
| J1613.8+3410 | OS 319 | 16 13 41.06 | +34 12 47.91 | fsrq | 1.398 | 132 | 4.34 (0.72) | 118 | 2.63 (0.79) |
| J1635.2+3809 | 4C +38.41 | 16 35 15.49 | +38 08 04.50 | FSRQ | 1.814 | 120 | 2.53 (0.35) | 105 | 2.81 (0.40) |
| J1637.7+4715 | 4C +47.44 | 16 37 45.13 | +47 17 33.83 | fsrq | 0.735 | 91 | 0.89 (0.07) | 34 | 1.03 (0.18) |
| J1637.9+5719 | OS 562 | 16 38 13.46 | +57 20 23.98 | fsrq | 0.751 | 144 | 1.45 (0.14) | 127 | 1.62 (0.44) |
| J1640.6+3945 | NRAO 512 | 16 40 29.63 | +39 46 46.03 | FSRQ | 1.672 | 91 | 0.98 (0.15) | 30 | 0.81 (0.20) |
| J1642.9+3950 | 3C 345 | 16 42 58.81 | +39 48 36.99 | FSRQ | 0.593 | 313 | 5.50 (0.45) | 437 | 5.91 (0.72) |
| J1653.9+3945 | Mkn 501 | 16 53 52.22 | +39 45 36.61 | BLL | 0.033 | 113 | 1.47 (0.09) | 41 | 1.05 (0.30) |
| J1715.7+6837 | S4 1716+68 | 17 16 13.94 | +68 36 38.74 | fsrq | 0.777 | 55 | 0.55 (0.05) | … | … |
| J1727.1+4531 | S4 1726+45 | 17 27 27.65 | +45 30 39.73 | fsrq | 0.717 | 132 | 1.12 (0.32) | 49 | 1.03 (0.27) |
| J1734.3+3858 | B2 1732+38A | 17 34 20.58 | +38 57 51.44 | fsrq | 0.976 | 215 | 0.97 (0.20) | 251 | 1.49 (0.56) |
| J1740.3+4736 | S4 1738+47 | 17 39 57.13 | +47 37 58.36 | fsrq | … | 95 | 0.81 (0.09) | 25 | 0.79 (0.28) |
| J1740.3+5211 | 4C +51.37 | 17 40 36.98 | +52 11 43.41 | fsrq | 1.381 | 123 | 1.11 (0.26) | 42 | 1.24 (0.55) |
| J1748.6+7005 | S4 1749+70 | 17 48 32.84 | +70 05 50.77 | bll | 0.770 | 136 | 0.82 (0.21) | 44 | 0.85 (0.33) |
| J1751.5+0939 | OT 081 | 17 51 32.82 | +09 39 00.73 | bll | 0.322 | 112 | 2.02 (0.46) | 121 | 2.87 (0.64) |
| J1800.5+7827 | S5 1803+784 | 18 00 45.68 | +78 28 04.02 | bll | 0.691 | 170 | 2.54 (0.21) | 183 | 2.33 (0.24) |
| J1801.5+4403 | S4 1800+44 | 18 01 32.31 | +44 04 21.90 | fsrq | 0.663 | 103 | 0.95 (0.20) | 32 | 1.14 (0.40) |
| J1806.7+6949 | 3C 371 | 18 06 50.68 | +69 49 28.11 | fsrq | 0.049 | 184 | 2.06 (0.25) | 266 | 2.02 (0.42) |
| J1824.2+5649 | 4C +56.27 | 18 24 07.07 | +56 51 01.49 | bll | 0.664 | 141 | 1.25 (0.13) | 131 | 1.26 (0.21) |
| J1829.6+4844 | 3C 380 | 18 29 31.78 | +48 44 46.16 | css | 0.692 | 132 | 5.26 (0.14) | 127 | 2.20 (0.28) |
| J1849.2+6705 | S4 1849+67 | 18 49 16.07 | +67 05 41.68 | FSRQ | 0.657 | 136 | 1.05 (0.21) | 124 | 1.70 (0.41) |
| J1927.7+6118 | S4 1926+61 | 19 27 30.44 | +61 17 32.88 | bll | 156 | 1.14 (0.23) | 131 | 1.20 (0.38) | |
| J2005.2+7752 | S5 2007+77 | 20 05 30.93 | +77 52 43.14 | bll | 0.342 | 146 | 1.04 (0.18) | 68 | 1.03 (0.25) |
| J2007.3+6605 | TXS 2007+659 | 20 07 28.77 | +66 07 22.54 | fsrq | 1.325 | 133 | 0.86 (0.21) | 39 | 0.74 (0.17) |
| J2010.3+7228 | 4C +72.28 | 20 09 52.30 | +72 29 19.35 | bll | … | 161 | 1.10 (0.08) | 117 | 0.96 (0.15) |
| J2022.5+7612 | S5 2023+760 | 20 22 35.58 | +76 11 26.17 | bll | 0.590 | 76 | 0.67 (0.07) | 40 | 0.72 (0.11) |
| J2023.2+3154 | 4C +31.56 | 20 23 19.02 | +31 53 02.31 | bcu | 0.356 | 136 | 1.48 (0.15) | 36 | 1.13 (0.21) |
| J2025.2+3340 | B2 2023+33 | 20 25 10.84 | +33 43 00.21 | bll | 0.219 | 138 | 2.16 (0.42) | 120 | 2.88 (0.77) |
| J2031.8+1223 | PKS 2029+121 | 20 31 54.99 | +12 19 41.34 | bll | 1.215 | 238 | 1.19 (0.12) | 50 | 0.88 (0.13) |
| J2038.8+5113 | 3C 418 | 20 38 37.03 | +51 19 12.66 | fsrq | 1.686 | 135 | 4.28 (0.26) | 73 | 3.66 (0.84) |
| J2110.3+3540 | B2 2107+35A | 21 09 31.88 | +35 32 57.60 | bcu | 0.202 | 136 | 2.25 (0.60) | 112 | 2.49 (0.41) |
| J2123.6+0533 | OX 036 | 21 23 44.52 | +05 35 22.09 | fsrq | 1.940 | 165 | 2.04 (0.34) | 96 | 1.99 (0.46) |
| J2201.7+5047 | NRAO 676 | 22 01 43.54 | +50 48 56.39 | fsrq | 1.899 | 102 | 0.79 (0.07) | … | … |
| J2202.7+4217 | BL Lacertae | 22 02 43.29 | +42 16 39.98 | BLL | 0.067 | 457 | 2.56 (0.86) | 371 | 5.45 (2.44) |
| J2203.7+3143 | 4C +31.63 | 22 03 14.98 | +31 45 38.27 | FSRQ | 0.295 | 69 | 2.30 (0.34) | … | … |
| J2212.0+2355 | PKS 2209+236 | 22 12 05.97 | +23 55 40.54 | fsrq | 1.127 | 42 | 0.80 (0.17) | 21 | 0.59 (0.11) |
| J2232.5+1143 | CTA 102 | 22 32 36.41 | +11 43 50.90 | FSRQ | 1.032 | 318 | 4.22 (0.33) | 362 | 4.06 (1.11) |
| J2236.3+2829 | B2 2234+28A | 22 36 22.47 | +28 28 57.41 | fsrq | 0.790 | 144 | 1.57 (0.20) | 109 | 1.78 (0.41) |
| J2254.0+1608 | 3C 454.3 | 22 53 57.75 | +16 08 53.56 | FSRQ | 0.859 | 336 | 17.08 (2.39) | 448 | 11.69 (3.27) |
| J2321.9+2732 | 4C +27.50 | 23 21 59.86 | +27 32 46.44 | fsrq | 1.255 | 138 | 1.23 (0.12) | 103 | 1.44 (0.59) |
| J2330.5+1104 | 4C +10.73 | 23 30 40.85 | +11 00 18.71 | fsrq | 1.501 | 103 | 1.04 (0.08) | 24 | 0.74 (0.12) |
| J2354.1+4605 | 4C +45.51 | 23 54 21.68 | +45 53 04.24 | fsrq | 1.992 | 97 | 1.11 (0.08) | … | … |
| SMMAN Source | UMRAO | F-GAMMA | RATAN-600 | OVRO | ROBIN | MRO | SMA |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
| TXS 0008+704 | … | … | … | … | … | … | … |
| 4C +60.01 | … | … | … | … | … | … | 0014+612 |
| PKS 0056-00 | … | … | 5BZQJ0059+0006 | … | … | … | … |
| TXS 0059+581 | 0059+581 | J0102+5824 | 5BZUJ0102+5824 | 0059+581 | … | 0059+581 | 0102+584 |
| 4C +01.02 | 0106+013 | J0108+0135 | 5BZQJ0108+0135 | J0108+0135 | … | 0106+013 | 0108+015 |
| 4C +67.04 | … | … | 5BZUJ0110+6805 | TB0110+6805 | … | … | … |
| 4C +14.06 | … | … | … | … | … | … | … |
| OC 457 | 0133+476 | J0136+4751 | 5BZQJ0136+4751 | J0136+4751 | … | 0133+476 | … |
| 4C +15.05 | 0202+149 | … | 5BZUJ0204+1514 | J0204+1514 | … | … | 0204+152 |
| S5 0212+73 | 0212+735 | J0217+734 | 5BZQJ0217+7349 | J0217+7349 | … | 0212+735 | 0217+738 |
| B2 0218+357 | 0218+357 | J0221+3556 | 5BZUJ0221+3556 | J0221+3556 | … | 0218+357 | … |
| 3C 66A | … | J0222+4302 | 5BZBJ0222+4302 | 3C 66A | 0219+428 | 3C 66A | … |
| 4C +28.07 | 0234+285 | J0237+2848 | 5BZQJ0237+2848 | J0237+2848 | … | 0234+285 | 0237+288 |
| AO 0235+164 | 0235+164 | J0238+1636 | 5BZBJ0238+1636 | J0238+1636 | 0235+164 | 0235+164 | 0238+166 |
| OD 166 | … | … | … | J0242+1101 | … | … | 0242+110 |
| 4C +47.08 | 0300+470 | J0303+4716 | 5BZBJ0303+4716 | 0300+471 | … | … | 0303+472 |
| NGC 1218 | … | … | … | NGC1218 | … | … | … |
| NGC 1275 | 0316+413 | J0319+4130 | 5BZUJ0319+4130 | J0319+4130 | 0316+413 | 3C 84 | 0319+415 |
| NRAO 140 | 0333+321 | J0336+3218 | 5BZQJ0336+3218 | 0333+321 | … | 0333+321 | 0336+323 |
| B3 0350+465 | … | … | … | … | … | … | 0354+467 |
| TXS 0354+599 | … | … | … | TXS0354+599 | … | … | … |
| 3C 111 | 0415+379 | J0418+3801 | … | 0415+379 | … | 0415+379 | 0418+380 |
| 4C +41.11 | … | … | 5BZBJ0423+4150 | TB0423+4150 | … | … | 0423+418 |
| PKS 0507+17 | … | J0510+1800 | 5BZQJ0510+1800 | J0510+1800 | … | 0507+179 | 0510+180 |
| B3 0509+406 | … | … | … | B30509+406 | … | … | … |
| 4C +45.08 | … | … | … | 4C+45.08 | … | … | … |
| 4C +27.15 | … | … | … | … | … | … | … |
| TXS 0518+211 | … | J0521+2112 | 5BZBJ0521+2112 | … | … | RGBJ0521+212 | … |
| OG 050 | 0529+075 | J0532+0732 | 5BZQJ0532+0732 | J0532+0732 | … | … | 0532+075 |
| TXS 0529+483 | … | … | 5BZQJ0533+4822 | TXS0529+483 | … | … | 0533+483 |
| 4C +22.12 | … | … | … | … | … | … | … |
| S5 0633+73 | … | … | 5BZQJ0639+7324 | J0639+7324 | … | … | … |
| S4 0707+47 | 0707+476 | … | 5BZBJ0710+4732 | B30707+476 | … | 0707+476 | 0710+475 |
| S5 0716+71 | 0716+714 | J0721+7120 | 5BZBJ0721+7120 | J0721+7120 | 0716+714 | 0716+714 | 0721+713 |
| OI 280 | 0748+126 | J0750+1231 | 5BZQJ0750+1231 | J0750+1231 | … | 0748+126 | 0750+125 |
| PKS 0754+100 | 0754+100 | … | 5BZBJ0757+0956 | J0757+0956 | … | 0754+100 | 0757+099 |
| NGC 2484 | 0755+379 | … | … | … | … | … | … |
| TXS 0800+618 | … | … | … | J0805+6144 | … | … | … |
| S4 0805+41 | 0805+410 | … | 5BZQJ0808+4052 | J0808+4052 | … | … | … |
| S4 0814+42 | 0814+425 | J0818+4222 | … | J0818+4222 | … | 0814+425 | 0818+423 |
| OJ 535 | 0820+560 | J0824+5552 | 5BZQJ0824+5552 | J0824+5552 | … | 0820+560 | 0824+558 |
| 4C +39.23 | 0821+394 | … | 5BZQJ0824+3916 | J0824+3916 | … | 0821+394 | 0824+392 |
| OJ 248 | 0827+243 | J0830+2410 | 5BZQJ0830+2410 | J0830+2410 | 0827+243 | OJ 248 | 0830+241 |
| OJ 448 | 0828+493 | … | 5BZBJ0832+4913 | … | … | 0828+493 | … |
| 3C 207 | 0838+133 | … | 5BZUJ0840+1312 | 3C 207 | 0838+133 | … | 3C207 |
| 0836+710 | 0836+710 | J0841+7053 | … | 0836+710 | 0836+710 | 0836+710 | … |
| OJ 287 | 0851+202 | J0854+2006 | 5BZBJ0854+2006 | J0854+2006 | 0851+202 | OJ 287 | 0854+201 |
| S4 0859+47 | 0859+470 | … | 5BZQJ0903+4651 | S4 0859+47 | … | 0859+470 | 0903+468 |
| SMMAN Sources | UMRAO | F-GAMMA | RATAN-600 | OVRO | ROBIN | MRO | SMA |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
| S4 0900+42 | … | … | … | S4 0900+42 | … | … | … |
| PKS 0906+01 | 0906+015 | … | … | J0909+0121 | … | 0909+013 | 0909+013 |
| S4 0913+39 | … | … | 5BZQJ0916+3854 | S40913+39 | … | … | … |
| S4 0917+44 | 0917+449 | J0920+4441 | 5BZQJ0920+4441 | J0920+4441 | … | 0917+449 | 0920+446 |
| OK 630 | 0917+624 | … | 5BZQJ0921+6215 | J0921+6215 | … | 0921+622 | 0921+622 |
| 4C +40.24 | 0945+408 | … | … | J0948+4039 | … | 0945+408 | 0948+406 |
| 4C +55.17 | 0954+556 | … | … | C0957+5522 | … | S4 0954+556 | 0957+553 |
| OK 492 | 0955+476 | … | 5BZQJ0958+4725 | J0958+4725 | … | 0958+474 | 0958+474 |
| S4 0954+65 | 0954+658 | J0958+6533 | 5BZBJ0958+6533 | J0958+6533 | 0954+658 | S4 0954+65 | 0958+655 |
| 4C +40.25 | 1020+400 | … | 5BZQJ1023+3948 | J1023+3948 | … | … | … |
| S4 1030+41 | 1030+415 | … | 5BZQJ1033+4116 | J1033+4116 | … | 1030+415 | … |
| S4 1030+61 | … | J1033+6051 | 5BZQJ1033+6051 | J1033+6051 | … | … | 1033+608 |
| B2 1040+24 | … | … | 5BZQJ1043+2408 | J1043+2408 | … | TEX 1040+244 | 1043+241 |
| S5 1039+81 | … | … | … | J1044+8054 | … | 1039+811 | 1044+809 |
| S5 1044+71 | … | … | 5BZUJ1048+7143 | J1048+7143 | … | … | 1048+717 |
| 4C +01.28 | 1055+018 | J1058+0133 | 5BZUJ1058+0133 | J1058+0133 | … | 1055+018 | 1058+015 |
| TXS 1125+596 | … | J1128+5925 | … | J1128+5925 | … | … | … |
| B2 1128+38 | 1128+385 | … | 5BZQJ1130+3815 | J1130+3815 | … | … | … |
| 4C +10.33 | … | … | … | … | … | … | … |
| 3C 264 | … | … | … | 3C 264 | … | … | … |
| S4 1144+40 | 1144+402 | … | 5BZQJ1146+3958 | J1146+3958 | … | … | 1146+399 |
| OM 484 | 1150+497 | … | 5BZQJ1153+4931 | OM484 | … | 1150+497 | 1153+495 |
| B3 1151+408 | … | … | … | … | … | … | … |
| Ton 599 | 1156+295 | J1159+2914 | 5BZQJ1159+2914 | J1159+2914 | 1156+295 | 4C 29.45 | 1159+292 |
| 4C +13.46 | … | … | 5BZQJ1213+1307 | 4C +13.46 | … | … | … |
| 4C +21.35 | 1222+216 | J1224+2122 | 5BZQJ1224+2122 | C1224+2122 | 1222+21 | PKS 1222+216 | 1224+213 |
| 3C 273 | 1226+023 | J1229+0203 | 5BZQJ1229+0203 | J1229+0203 | 1226+023 | 3C 273 | 1229+020 |
| M 87 | … | J1230+1223 | … | 3C 274 | … | 3C 274 | 1230+123 |
| PKS 1236+077 | 1236+077 | … | 5BZUJ1239+0730 | J1239+0730 | … | … | 1239+075 |
| 3C 279 | 1256-057 | J1256-0547 | 5BZQJ1256-0547 | J1256-0547 | 1253–055 | 3C 279 | 1256-057 |
| ON 393 | … | … | … | J1257+3229 | … | … | … |
| TXS 1300+580 | … | … | 5BZUJ1302+5748 | J1302+5748 | … | 1300+580 | … |
| 4C +12.46 | … | … | 5BZBJ1309+1154 | J1309+1154 | … | MC 21307+12 | 1309+119 |
| OP 313 | 1308+326 | J1330+3220 | 5BZUJ1310+3220 | J1310+3220 | … | 1308+326 | 1310+323 |
| PKS 1348+007 | … | … | 5BZQJ1351+0031 | PKS 1348+007 | … | … | … |
| OQ 530 | 1418+546 | … | 5BZBJ1419+5423 | J1419+5423 | … | OQ 530 | 1419+543 |
| 3C 303 | 1441+522 | … | … | … | … | … | … |
| TXS 1459+480 | … | … | 5BZQJ1500+4751 | J1500+4751 | … | … | … |
| 4C +14.60 | 1538+149 | … | 5BZBJ1540+1447 | J1540+1447 | … | 4C +14.60 | 1540+147 |
| PKS 1546+027 | 1546+027 | … | 5BZQJ1549+0237 | J1549+0237 | … | 1546+027 | 1549+026 |
| 4C +05.64 | 1548+056 | J1550+0527 | 5BZQJ1550+0527 | J1550+0527 | … | 1548+056 | 1550+054 |
| GB6 J1604+5714 | … | … | … | J1604+5714 | … | … | … |
| OS 319 | 1611+343 | J1613+3412 | 5BZQJ1613+3412 | J1613+3412 | 1611+343 | DA 406 | 1550+054 |
| 4C +38.41 | 1633+382 | J1635+3808 | 5BZQJ1635+3808 | J1635+3808 | 1633+382 | 4C +38.41 | 1635+381 |
| 4C +47.44 | … | … | 5BZQJ1637+4717 | J1637+4717 | … | … | 1637+472 |
| OS 562 | 1637+574 | J1638+5720 | … | J1638+5720 | … | 1637+574 | 1638+573 |
| NRAO 512 | 1638+398 | … | 5BZQJ1640+3946 | J1640+3946 | … | 1638+398 | 1640+397 |
| SMMAN Sources | UMRAO | F-GAMMA | RATAN-600 | OVRO | ROBIN | MRO | SMA |
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
| 3C 345 | 1641+399 | J1642+3948 | 5BZQJ1642+3948 | J1642+3948 | 1641+399 | 3C 345 | 1642+398 |
| Mkn 501 | 1652+398 | J1653+3945 | 5BZBJ1653+3945 | J1653+3945 | 1652+398 | MARK 501 | 1653+397 |
| S4 1716+68 | … | … | … | J1716+6836 | … | … | 1716+686 |
| S4 1726+45 | 1726+455 | … | 5BZQJ1727+4530 | J1727+4530 | … | … | 1727+455 |
| B2 1732+38A | 1732+389 | … | … | J1734+3857 | … | … | 1734+389 |
| S4 1738+47 | 1738+476 | … | 5BZBJ1739+4737 | J1739+4737 | … | … | … |
| 4C +51.37 | 1739+522 | … | 5BZQJ1740+5211 | J1740+5211 | … | S4 1739+52 | 1740+521 |
| S4 1749+70 | 1749+701 | J1748+7005 | 5BZBJ1748+7005 | J1748+7005 | … | … | … |
| OT 081 | 1749+096 | J1751+0939 | 5BZBJ1751+0939 | J1751+0939 | … | PKS 1749+096 | 1751+096 |
| S5 1803+784 | 1803+784 | J1800+7828 | 5BZBJ1800+7828 | J1800+7828 | … | S5 1803+784 | 1800+784 |
| S4 1800+44 | 1800+440 | … | 5BZQJ1801+4404 | J1801+4404 | … | … | 1801+440 |
| 3C 371 | 1807+698 | J1806+6949 | 5BZBJ1806+6949 | J1806+6949 | 1807+698 | 3C 371 | 1806+698 |
| 4C +56.27 | 1823+568 | J1824+5651 | 5BZBJ1824+5651 | J1824+5651 | … | 4C 56.27 | 1824+568 |
| 3C 380 | 1828+487 | … | 5BZUJ1829+4844 | 3C 380 | 1828+487 | 1828+487 | 1829+487 |
| S4 1849+67 | … | J1849+6705 | … | J1849+6705 | … | J1849+6705 | 1849+670 |
| S4 1926+61 | 1926+611 | … | 5BZBJ1927+6117 | J1927+6117 | … | … | 1927+612 |
| S5 2007+77 | 2007+777 | … | … | J2005+7752 | … | S52007+77 | 2005+778 |
| TXS 2007+659 | … | … | … | J2007+6607 | … | … | … |
| 4C +72.28 | 2010+723 | … | 5BZBJ2009+7229 | J2009+7229 | … | … | 2009+724 |
| S5 2023+760 | … | … | 5BZBJ2022+7611 | J2022+7611 | … | … | … |
| 4C +31.56 | … | … | … | 2021+317 | … | … | 2023+318 |
| B2 2023+33 | … | … | … | 3EGJ2027+3429 | … | … | 2025+337 |
| PKS 2029+121 | 2029+121 | J2031+1219 | 5BZUJ2031+1219 | J2031+1219 | … | … | … |
| 3C 418 | 2037+511 | … | … | 2037+511 | … | 2037+511 | 2038+513 |
| B2 2107+35A | … | … | … | … | … | … | 2109+355 |
| OX 036 | 2121+053 | … | 5BZQJ2123+0535 | J2123+0535 | … | 2121+053 | 2123+055 |
| NRAO 676 | … | … | … | NRAO 676 | … | … | … |
| BL Lacertae | 2200+420 | J2202+4216 | 5BZBJ2202+4216 | BL Lacertae | 2200+420 | BL LAC | 2202+422 |
| 4C +31.63 | 2201+315 | J2203+3145 | 5BZQJ2203+3145 | J2203+3145 | … | 2201+315 | 2203+317 |
| PKS 2209+236 | … | J2212+2355 | 5BZUJ2212+2355 | J2212+2355 | … | … | … |
| CTA 102 | 2230+114 | J2232+1143 | 5BZQJ2232+1143 | 2230+114 | 2230+114 | 2230+114 | 2232+117 |
| B2 2234+28A | 2234+282 | … | 5BZQJ2236+2828 | J2236+2828 | … | 2234+282 | 2236+284 |
| 3C 454.3 | 2251+158 | J2253+1608 | 5BZQJ2253+1608 | J2253+1608 | 2251+158 | 3C 454.3 | 2253+161 |
| 4C +27.50 | 2319+272 | … | 5BZQJ2321+2732 | J2321+2732 | … | … | 2321+275 |
| 4C +10.73 | 2328+107 | … | 5BZQJ2330+1100 | J2330+1100 | … | … | … |
| 4C +45.51 | 2351+456 | … | 5BZQJ2354+4553 | 2351+456 | … | … | 2354+458 |
| SMMAN source | Flux (0.1–100 GeV) | LAT | (0.1–100 GeV) | |||||||
| 4.8 GHz | 23.6 GHz | 4.8 GHz | 23.6 GHz | W Hz-1 | W Hz-1 | () | (photon index) | () | ||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | (11) |
| TXS 0008+704 | … | … | … | … | … | |||||
| 4C +60.01 | … | … | … | … | … | |||||
| PKS 0056-00 | … | … | … | … | … | … | … | … | ||
| TXS 0059+581 | 3.80 | |||||||||
| 4C +01.02 | ||||||||||
| 4C +67.04 | … | … | … | … | … | … | … | … | ||
| 4C +14.06 | … | … | … | … | … | … | … | … | ||
| OC 457 | ||||||||||
| 4C +15.05 | ||||||||||
| S5 0212+73 | ||||||||||
| B2 0218+357 | ||||||||||
| 3C 66A | … | … | … | … | … | |||||
| 4C +28.07 | ||||||||||
| AO 0235+164 | ||||||||||
| OD 166 | ||||||||||
| 4C +47.08 | ||||||||||
| NGC 1218 | ||||||||||
| NGC 1275 | ||||||||||
| NRAO 140 | ||||||||||
| B3 0350+465 | … | … | … | |||||||
| TXS 0354+599 | ||||||||||
| 3C 111 | ||||||||||
| 4C +41.11 | … | … | … | |||||||
| PKS 0507+17 | ||||||||||
| B3 0509+406 | … | … | … | |||||||
| 4C +45.08 | … | … | … | … | … | … | … | … | ||
| 4C +27.15 | … | … | … | … | … | … | … | … | ||
| TXS 0518+211 | … | … | … | … | … | |||||
| OG 050 | ||||||||||
| TXS 0529+483 | ||||||||||
| 4C +22.12 | … | … | … | |||||||
| S5 0633+73 | ||||||||||
| S4 0707+47 | … | … | … | … | … | |||||
| S5 0716+71 | ||||||||||
| OI 280 | ||||||||||
| PKS 0754+100 | ||||||||||
| NGC 2484 | … | … | … | … | … | … | … | … | ||
| TXS 0800+618 | … | … | … | … | … | |||||
| S4 0805+41 | ||||||||||
| S4 0814+42 | … | … | … | … | … | |||||
| OJ 535 | … | … | … | … | ||||||
| 4C +39.23 | ||||||||||
| OJ 248 | ||||||||||
| OJ 448 | … | … | … | … | … | … | … | … | ||
| 3C 207 | ||||||||||
| 0836+710 | ||||||||||
| OJ 287 | ||||||||||
| S4 0859+47 | … | … | … | |||||||
| SMMAN source | Flux (0.1–100 GeV) | LAT | (0.1–100 GeV) | |||||||
| 4.8 GHz | 23.6 GHz | 4.8 GHz | 23.6 GHz | W Hz-1 | W Hz-1 | () | (photon index) | () | ||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | (11) |
| S4 0900+42 | … | … | … | … | … | … | … | … | ||
| PKS 0906+01 | ||||||||||
| S4 0913+39 | … | … | … | … | … | |||||
| S4 0917+44 | ||||||||||
| OK 630 | ||||||||||
| 4C +40.24 | … | … | … | … | … | |||||
| 4C +55.17 | ||||||||||
| OK 492 | ||||||||||
| S4 0954+65 | ||||||||||
| 4C +40.25 | … | … | … | … | … | |||||
| S4 1030+41 | ||||||||||
| S4 1030+61 | … | … | … | |||||||
| B2 1040+24 | ||||||||||
| S5 1039+81 | ||||||||||
| S5 1044+71 | ||||||||||
| 4C +01.28 | ||||||||||
| TXS 1125+596 | ||||||||||
| B2 1128+38 | ||||||||||
| 4C +10.33 | … | … | … | … | … | … | … | … | ||
| 3C 264 | … | … | … | … | … | … | … | … | ||
| S4 1144+40 | ||||||||||
| OM 484 | ||||||||||
| B3 1151+408 | … | … | … | … | … | |||||
| Ton 599 | ||||||||||
| 4C +13.46 | … | … | … | … | … | … | … | … | ||
| 4C +21.35 | ||||||||||
| 3C 273 | ||||||||||
| M 87 | ||||||||||
| PKS 1236+077 | … | … | … | … | …. | … | … | … | ||
| 3C 279 | ||||||||||
| ON 393 | … | … | … | … | … | |||||
| TXS 1300+580 | … | … | … | … | … | |||||
| 4C +12.46 | … | … | … | |||||||
| OP 313 | ||||||||||
| PKS 1348+007 | … | … | … | … | … | |||||
| OQ 530 | ||||||||||
| 3C 303 | … | … | … | |||||||
| TXS 1459+480 | … | … | … | … | … | … | … | … | ||
| 4C +14.60 | … | … | … | … | … | … | … | … | ||
| PKS 1546+027 | … | … | … | … | … | |||||
| 4C +05.64 | … | … | … | … | … | |||||
| GB6 J1604+5714 | … | … | … | … | … | |||||
| OS 319 | ||||||||||
| 4C +38.41 | ||||||||||
| 4C +47.44 | ||||||||||
| OS 562 | ||||||||||
| NRAO 512 | ||||||||||
| SMMAN source | Flux (0.1–100 GeV) | LAT | (0.1–100 GeV) | |||||||
| 4.8 GHz | 23.6 GHz | 4.8 GHz | 23.6 GHz | W Hz-1 | W Hz-1 | () | (photon index) | () | ||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | (11) |
| 3C 345 | ||||||||||
| Mkn 501 | ||||||||||
| S4 1716+68 | … | … | … | … | … | |||||
| S4 1726+45 | ||||||||||
| B2 1732+38A | ||||||||||
| S4 1738+47 | … | … | … | |||||||
| 4C +51.37 | ||||||||||
| S4 1749+70 | ||||||||||
| OT 081 | ||||||||||
| S5 1803+784 | ||||||||||
| S4 1800+44 | ||||||||||
| 3C 371 | ||||||||||
| 4C +56.27 | ||||||||||
| 3C 380 | ||||||||||
| S4 1849+67 | ||||||||||
| S4 1926+61 | ||||||||||
| S5 2007+77 | ||||||||||
| TXS 2007+659 | ||||||||||
| 4C +72.28 | … | … | … | |||||||
| S5 2023+760 | ||||||||||
| 4C +31.56 | … | … | … | |||||||
| B2 2023+33 | ||||||||||
| PKS 2029+121 | ||||||||||
| 3C 418 | ||||||||||
| B2 2107+35A | … | … | … | |||||||
| OX 036 | ||||||||||
| NRAO 676 | … | … | … | … | … | |||||
| BL Lacertae | ||||||||||
| 4C +31.63 | … | … | … | … | … | … | … | … | ||
| PKS 2209+236 | … | … | … | |||||||
| CTA 102 | ||||||||||
| B2 2234+28A | ||||||||||
| 3C 454.3 | ||||||||||
| 4C +27.50 | … | … | … | |||||||
| 4C +10.73 | ||||||||||
| 4C +45.51 | … | … | … | … | … | … | … | … | ||
Appendix B Radio Light curves of Individual Sources Monitored in SMMAN Program
Figure 19 of Appendix B presents the light curves for all the 131 SMMAN sources and the three calibrator sources in the program at 4.8 and 23.6 GHz. The data points shown here are unbinned. For some sources, either 4.8 GHz or 23.6 GHz light curves are available; accordingly, those are presented.